# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v5

@@ Abstract
Continual learning (CL) aims to accumulate knowledge from sequential data and task streams. Leveraging their strong generalization and flexibility, pre-trained vision-language embedding models such as CLIP (Contrastive Language-Image Pre-training) have been widely adopted and validated in CL. In addition to learning new knowledge, we investigate whether the pre-trained knowledge in CLIP, can be retained or even enhanced, in CL, while incorporating new knowledge from a data stream. Existing CL methods primarily focus on continual downstream adaptation using components isolated from the pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge by relying on additional reference data, resulting in high training costs. To address these limitations, we propose a universal and efficient CL approach for CLIP based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL. By analyzing how LoRA rank and placement affect learning and forgetting in CL, we design CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data. This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability). Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities. Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead. Extensive experiments show that our approach enhances representations for new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
@@ 1 Introduction
Continual learning (CL) [18, 9, 49, 29, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [39, 36, 17] or the need to retrain on all previously seen data. With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context [70, 37, 48, 24, 58]. Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [44, 54, 24, 37]), with an emphasis on downstream task performance. While these task-specific designs [70, 37, 58] yield promising results, preserving the inherent capabilities of PTMs during CL is critical to maintaining and improving generalizable representations. Benefiting from the strong generalization and flexibility of jointly learned vision-text embeddings, vision-language models such as CLIP have been widely adopted and studied in CL [24, 66, 50, 69], including continual updating [17] or multi-domain CL [66] of CLIP.
Some approaches, such as ZSCL [66], aim to update PTM, i.e., CLIP, by incrementally incorporating new data through full fine-tuning while mitigating forgetting via memory replay [29, 54, 43] using additional reference data. However, this leads to high training costs and still struggles with forgetting (Fig. 1(a)). Other methods [61, 58] retain PTM capabilities during downstream CL by introducing task-specific components that are isolated from the PTM, requiring task/domain prediction or gating at inference. This increases inference complexity and limits adaptability to unseen data, restricting the model’s generalization potential (Fig. 1(b)).
@@
Figure 1: CL methods for CLIP: (a) Continual full fine-tuning uses reference data to apply universal updates and mitigate forgetting of pre-trained knowledge (e.g., zero-shot performance on unseen data [66]), but is computationally expensive and still susceptible to forgetting; (b) Task-specific modular methods (e.g., [61, 58]) introduce isolated components that require domain prediction or gating at inference, increasing complexity and limiting adaptability and generalization; (c) Our approach enables universal, parameter- and data-efficient adaptation without reference data or task-specific modules. It continually improves the pre-trained model and achieves superior performance on unseen data.
In this work, we propose a practical and efficient CL approach for CLIP that continually trains and updates pre-trained models without relying on downstream task-specific components. Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams. To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [21] for CL in PTMs. Rather than restricting LoRA to fixed locations as in prior works [30, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning. We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec. 3.2).
@@
Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates by dynamically optimizing for a minimized LoRA rank suitable for the current data. The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [64, 12, 32, 38, 63], with learnable importance weights corresponding to the singular values. With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training. This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting. Our contributions are fourfold:
  1. 1.
    We investigate a universal and practical CL paradigm for efficiently updating pre-trained CLIP on new data streams while mitigating forgetting, without reference data or additional components.
  2. 2.
    We conduct in-depth analyses on the impact of LoRA ranks and placements, revealing their crucial role in balancing downstream learning and knowledge retention.
  3. 3.
    We propose CoDyRA, which dynamically optimizes a minimized rank for parameter updates based on current data, with learnable importance weights for each rank.
  4. 4.
    Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, and additionally prove superiority on pre-trained knowledge retention using hold-out data.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [34, 3, 6, 31, 7, 59, 60] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [25, 1, 62, 2, 23] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [46, 47, 68, 48, 37, 30, 53, 52, 44, 51] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. Continual learning of CLIP [24, 65] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. [17] continually adapt CLIP to temporal distribution shifts via timestamped data updates. The approach introduced in ZSCL [66] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [61, 45, 58] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [58] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [21] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, [56] learns a small set of task-specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [63, 38], treating each rank independently [12, 64, 32], or employing a mixture of ranks [55]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
Figure 2: Overview of CoDyRA: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
@@ 3.1 Preliminaries
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multilayer perceptron (MLP). In each transformer layer, the attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
The MLP module within each transformer block consists of two linear transformations separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [20]. The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
@@
Vision-language CLIP model. In CLIP model with joint vision-language embedding [42, 22, 14, 28], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
@@
CL of CLIP. We focus on CL of CLIP with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [66, 61] and cross-domain task-agnostic incremental learning (X-TAIL) [58]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total. In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This poses a more challenging scenario, as the model must classify test images across a broader range of domains.

### v6

@@ Abstract
Continual learning (CL) aims to accumulate knowledge from sequential tasks without catastrophic forgetting. Vision–language models like CLIP, with strong generalization, are widely used for CL. Existing methods often adapt isolated PTM components, adding inference complexity and limiting PTM improvement, or rely on replay, stored information, or assumptions, incurring high costs and limited applicability. To advance models as continual learners, we explore CL via natural, efficient PTM updates instead of complex task-specific additions. We thus study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting. We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while a relatively lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation. Crucially, we find a plasticity–stability balance tied to rank across parameters and tasks, with moderately small ranks maximizing CL benefits. Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates PTMs with LoRA adapters for all parameters of adaptively optimized rank number. While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving a plasticity–stability balance tailored to different parameters and tasks. Despite allowing updates to all parameters, the adaptively selected and minimized LoRA ranks help keep the updated model closer to its previous state, while still enabling comprehensive and effective learning of new tasks. CoDyRA enables efficient CL as a sequence of LoRA-based tasks without storing past data, task information, or relying on assumptions. It preserves the original model architecture and deployment pipeline, adding no inference overhead. Extensive experiments show CoDyRA improves new representations while retaining old knowledge, achieving state-of-the-art results. Code is available.
@@ 1 Introduction
Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Wang et al., 2024b) focuses on incrementally learning from new data streams without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need to retrain on all previously seen data. With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context (Zhou et al., 2024; McDonnell et al., 2024; Wang et al., 2024a; Jha et al., 2024). Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL (Smith et al., 2023; Wortsman et al., 2022)), with an emphasis on downstream tasks. While task-specific and downstream-oriented designs (Zhou et al., 2024; McDonnell et al., 2024; Xu et al., 2024; Yu et al., 2024) show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement. To advance toward models that function as continual learners, we consider CL with PTMs that enable flexible and efficient updates, preserving or even enhancing generalizable representations while acquiring new knowledge.
Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL (Jha et al., 2024; Zheng et al., 2023; Wang et al., 2023; Zhou et al., 2023), including continual updating (Garg et al., 2023) and multi-domain CL (Zheng et al., 2023). Beyond downstream performance, the retention of knowledge in PTMs is also evaluated. Some approaches, such as ZSCL (Zheng et al., 2023), aim to continually update PTM, i.e., CLIP, through full fine-tuning while mitigating forgetting via memory replay (Li & Hoiem, 2017; Wortsman et al., 2022; Rebuffi et al., 2017b) using additional reference data. However, this leads to high training costs and still struggles with catastrophic forgetting (Fig. 1(a)). To directly control the forgetting, some methods (Yu et al., 2024; Xu et al., 2024) introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating at inference. These methods can deliver strong results with task-specific gating and branching; however, by isolating tasks and relying heavily on per-task modules and accurate task prediction, they become restricted to certain domains and limited in both general CL applicability and inference efficiency. Moreover, with restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig. 1(b)).
@@
Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance (Zheng et al., 2023)) but remains costly and prone to forgetting. (b) Task-specific modular methods (e.g., (Yu et al., 2024; Xu et al., 2024)) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization. (c) Our approach enables efficient, universal continual PTM updating without reference data or task-specific designs, retaining PTM knowledge and improving performance on unseen data.
Straightforward adaptation of PTMs can be achieved through direct model updating through full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods (Zhou et al., 2022c; b), such as Low-Rank Adaptation (LoRA) (Hu et al., 2021). This raises a fundamental question: instead of relying on complex additions or task-specific components, can CL be performed by leveraging these natural PTM update mechanisms? Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing extensive assumptions. We propose an efficient CL approach for CLIP that enables the model to be updated naturally across a stream of tasks, while retaining foundational pre-trained knowledge. We build upon LoRA for efficiency. While there have been some LoRA-based CL methods (Liang & Li, 2024; Wu et al., 2025; Zhu et al., 2025), they rely on specific regularization (e.g., orthogonal property) for specific settings.
@@
We first systematically analyze and ablate how LoRA adapters at different locations in the network with different ranks influence learning and forgetting (Sec. 3.2). Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Our analyses empirically demonstrate that, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks (i.e., plasticity) but tends to increase forgetting, while lower-rank LoRA mitigates forgetting (i.e., stability) but limits adaptation. Importantly, we identify that there exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates the PTM using LoRA adapters with dynamically and adaptively optimized ranks. CoDyRA performs new task learning while jointly minimizing the rank of each LoRA adaptively. Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank during training. By updating with LoRA adapters of minimized ranks, CoDyRA inherently biases models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
Our contributions are fourfold: (1) We study PTM-based CL with natural and efficient continual update of the model, not relying on complex task-specific additions, without storing past data, task information, or relying on assumptions. (2) We comprehensively analyze learning–forgetting trade-offs across LoRA ranks and placements, identifying a plasticity–stability balance at moderately small ranks and deriving design strategies. This study guides our CL approach with regularized forgetting and provides evidence for future research. (3) We propose CoDyRA, which adaptively minimizes LoRA ranks via sparsity-promoting regularization. It handles different parameters and tasks, yielding regularized updates that reduce interference and keep the model closer to its previous state with less forgetting. (4) Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, achieving a balanced trade-off between plasticity and stability.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks (Wang et al., 2022a; b; Zhou et al., 2022a; Wang et al., 2024a; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c) adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. Continual learning of CLIP (Jha et al., 2024; Zhang et al., 2024b) aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. (Garg et al., 2023) continually adapt CLIP to temporal distribution shifts via timestamped data updates. The approach introduced in ZSCL (Zheng et al., 2023) addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Huang et al., 2025) have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, (Xu et al., 2024) proposed a more challenging setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) (Hu et al., 2021) has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, (Wu et al., 2024b) learns a small set of task‐specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights (Zhang et al., 2024a; Meng et al., 2024), treating each rank independently (Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024), or employing a mixture of subspaces (Wu et al., 2024a). These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
@@ 3.1 Preliminaries
Transformers. A typical transformer model is composed of stacked blocks, each containing two main submodules: an multi-head attention (MHA) module (Attn) and a multilayer perceptron (MLP). The attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l}.$ The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively. The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU (Hendrycks & Gimpel, 2016). The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension.
Figure 2: Overview of CoDyRA: we propose dynamic rank-selection LoRA, enabling each pre-trained weight matrix to adaptively add necessary ranks for downstream adaptation while retaining pre-trained capabilities. After each task, dynamic rank updates are merged into the pre-trained weights with no inference overhead.
@@
CLIP model. In CLIP (Radford et al., 2021; Ilharco et al., 2021; Fang et al., 2024; Li et al., 2022), both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding. Given an image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, the vision encoder produces features ${\mathbf{z}}_{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. A class label $y$ is embedded into a prompt such as “a photo of a [CLS]” to form text input $\mathbf{t}$}, which the text encoder maps to ${\mathbf{z}}_{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. At inference, the probability of classifying $\mathbf{x}$} into class $y_{i}\in\{1,\dots,C\}$ is: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity and $\tau$ is a temperature parameter.
Figure 3: Task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion points and ranks. Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
Figure 4: Impact of removing trained LoRA modules on downstream task adaptation (Aircraft, left y-axis) and retention of existing capabilities (ImageNet-1k, right y-axis). “T” and “V” indicate text-encoder and vision-encoder modules, respectively.
@@
CL with CLIP. We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) (Zheng et al., 2023; Yu et al., 2024) and cross-domain task-agnostic incremental learning (X-TAIL) (Xu et al., 2024). In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$. In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.